solve (include all real and imaginary solutions)
t^4+t^2=20

Answer by stanbon(75887)

solve (include all real and imaginary solutions)
t^4+t^2=20
-----
t^4+t^2-20 = 0
---
(t^2+5)(t^2-4) = 0
---
t = +/-sqrt(5) or t = +/-2
==============================
Cheers,
Stan H.